# Towards Better Benchmark Datasets for Inductive Knowledge Graph Completion

### v2

@@ 3.3 Why is This So Common on Inductive Datasets?
The studies in this section show that the current strategy for constructing inductive datasets from existing transductive datasets is liable to introduce a well-performing shortcut into the existing graph. Currently, it almost always leads to a sharp increase in the PPR performance. Furthermore, attempting to limit the severity of this issue is very difficult while also generating train and inference graphs of reasonable sizes. This suggests that a new strategy is needed for sampling graphs for inductive KGC from transductive datasets.
@@
Figure 4: Experiments when generating inductive datasets from FB15k-237. (a) The generated datasets always have a much higher $\Delta\text{SPD}$ than the transductive dataset. (b)/(c) There is a trade-off between the size of the train and inference graph, making it difficult to obtain a good split.
4 Constructing Inductive Datasets through Graph Partitioning
@@
4.1 Partition-Based Dataset Sampling
@@
We note that the task of constructing inductive datasets from an existing graph $\mathcal{G}$ can be framed as a graph partitioning problem. Formally, we want to sample two non-overlapping partitions from the graph such that $\mathcal{G}_{\text{train}},\mathcal{G}_{\text{inf}}\subset\mathcal{G}$ and $\mathcal{G}_{\text{train}}\cap\mathcal{G}_{\text{inf}}=\emptyset$. Given the analysis in Section 3.2, we hope to sample subgraphs such that $\Delta\text{SPD}(\mathcal{G}_{\text{train}})\approx\Delta\text{SPD}(\mathcal{G}_{\text{inf}})\approx\Delta\text{SPD}(\mathcal{G}$). But, how do we find subgraphs that satisfy this property? Intuitively, we want to sample each subgraph so that it’s removal has little effect on the initial graph’s structure. We give an example in Figure 5 where we sample two subgraphs from an existing graph. As we can see, even though the graph is partitioned, the relationship between entities in the same partition remain roughly the same before and after the partition. This is because there already exists little relationship between the two partitions in the original graph.
Figure 5: Sampling Two Subgraphs
@@
Multiple popular approaches [29, 30] exist that attempt to divide the graph into optimal partitions. The guiding principle in these approaches is that the partitions should be internally dense but only sparsely connected to one another. Because of this, the entities in different partitions should only be weakly connected and have little impact on one another. Therefore, the relationship between entities in the same partition are minimally affected by outside entities or edges. As such, removing this partition from the graph should then have little effect on the entities in that partition. Also, since the partitions are created at the same time, we avoid sampling the inference graph after the training, which can be suboptimal.
@@
In practice, we consider using Spectral Clustering [29] or the Louvain method [30], as dependent on the dataset. Once the original graph is partitioned into $n$ partitions, we sample $k$ of those to be used. The partitions are chosen such that they display similar properties to the original graph (see Section 4.2 for more). Of the $k$ partitions, one is chosen as the training graph while the other $k-1$ are designated as separate inference graphs. This is an important advantage of using partitioning to sample the graphs, as it allows us to avoid having to sample multiple different train and inference pairs as in [7, 9], thereby allowing for more efficient benchmarking. Lastly, we note that when sampling graphs for the (E) inductive task, we further remove new relations from the inference graph. In practice, we find that we can easily find partitions where this amounts to little or no change in the inference graph. See Section B.2 for more details on the dataset creation process.
@@
Table 2: Mean $\Delta\text{SPD}$ and PPR Hits@10 of new and old Inductive (“Ind.”) datasets vs. their Transductive (“Trans.”) parent. We highlight the value of the inductive dataset closer to the transductive as blue. For every case, the $\Delta\text{SPD}$ and PPR Hits@10 of the new inductive datasets are more aligned with the transductive dataset than the older inductive datasets.
@@ 4.2 Analysis of New Datasets
In this section we analyze the new inductive datasets created following the partition-based procedure outlined in Section 4.1. We create (E) inductive datasets, from WN18RR [27], CoDEx-M [31], and HetioNet [32]. For the (E, R) datasets, we use CoDEx-M [31] and FB15k-237 [33]. Note that some datasets are only suitable for one task or another. For example, WN18RR and HetioNet have very few relations, making it nearly impossible to sample two partitions with significantly different relations for the (E, R) task. On the other hand, FB15k-237 contains too many relations to sample multiple graphs for the (E) tasks without removing many edges from either the train or inference graph.
@@
In Table 2 we show the PPR Hits@10 and $\Delta\text{SPD}$ for the inductive datasets and their original transductive dataset. When multiple inference graphs exist, we take the mean across each inference graph. When possible, we also include a comparison against those inductive datasets that already exist. For example, for WN18RR in the (E) task, 4 datasets exist from [7]. As compared to the old inductive datasets, the PPR performance for the new inductive datasets is much lower. We find that the average PPR performance is 78% lower on the new inductive datasets as compared to the older datasets. Furthermore, the PPR performance of the new inductive datasets is very similar to the performance on the original transductive dataset. A similar trend can be found when comparing the $\Delta\text{SPD}$. This analysis shows that newer sampling procedure can indeed sample inductive datasets that are much more similar to the original transductive graph, thereby greatly mitigating the PPR shortcut.
5 Experiments
5.1 Experimental Settings
@@
Datasets: We use the new datasets created in Section 4.1. For the (E) setting, this includes WN18RR, HetioNet, and CoDEx-M. For the (E, R) setting, it is FB15k-237 and CoDEx-M. We sample 2 inference graphs for each dataset, except for CoDEx-M on the (E) setting where we could only find 1 suitable graph for inference. For each inference graph, 10% of edges are randomly removed for testing. For validation 10% of edges are removed from the training graph. It is necessary that the validation samples are extracted from the train graph as the inference graphs must remain unobserved during training. The statistics for each dataset can be found in Appendix B.2.
@@
Baseline Methods: We consider prominent KGC methods including NBFNet [8], RED-GNN [11], NodePiece [19], InGram [9], and Neural LP [13]. We also consider the recent foundation model ULTRA [34]. We omit methods that have been shown to either be prohibitively expensive (e.g.,  [7, 20, 21]) or require the use of textual information [24, 25].
The full set of experimental settings can be found in Appendix D.
5.2 Results
@@
Main Results: The main results for the supervised methods can be found in Tables 3 and 4. We observe that on nearly every dataset, NBFNet and RED-GNN are the two best models. This indicates that conditional MPNNs [15], the class of model in which both belong to, are necessary for strong performance in inductive KGC. Interestingly, we observe that InGram struggles in the (E, R) setting. This is even true when the % of new relations is high. This runs counter to the results on older inductive datasets [9] where InGram excelled over NBFNet and RED-GNN. Lastly, we observe that Neural LP sometimes fails to converge, resulting in a near zero performance and thus causing the model to have a high performance variance.
Table 3: (E) Inductive Results (Hits@10) for supervised methods.
@@
Table 4: (E, R) Inductive Results for supervised methods. The % of new relations are in parentheses.
@@
Performance Comparison of PPR vs. SOTA: In our original analysis, we showed that for the older inductive datasets, the performance gap between the SOTA method of each dataset and PPR is quite small. Specifically, the mean % difference between the two was relatively small at 25-29% (see Figure 1). We now perform the same analysis on the new inductive datasets. The results are shown in Figure 6(a). We observe that PPR generally performs much worse than the SOTA method with a mean % difference of 99%. Furthermore, this is more in line with the transductive datasets which have a mean % difference of 121%. This suggests that in addition to the decrease in raw PPR performance, the relative performance of PPR also decreases on the newer datasets.
Table 5: Mean performance on Old vs. New inductive datasets by Method. Red indicates a decrease in the mean performance on the newer splits.
@@
Performance Comparison on New and Old Inductive Datasets: Previously, in Table 2, we compared both the PPR performance and $\Delta\text{SPD}$ on the new and older inductive datasets. We found that both metrics tend to be much higher on the older inductive datasets, indicating that our newer splits are effective in mitigating the shortcut. Given those results, a natural question is whether we see a similar drop in performance for neural methods? We limit our analysis to WN18RR (E) and FB15k-237 (E, R). This is either due to a lack of older datasets (i.e., CoDEX-M (E, R) and HetioNet (E)) or minimal results on the older datasets, i.e., CoDEX-M (E). For each dataset, we compute the mean performance across inference graphs. The results are shown in Table 5. We find that performance of most methods drops significantly across both datasets, with an average drop of 40.6% and 9.5% on WN18RR (E) and FB15k-237 (E, R), respectively. This suggests that removing the shortcut has a large negative effect on the performance, thereby suggesting that our new datasets are indeed harder.
@@
Performance of ULTRA [34]: We further compare against ULTRA, a recent foundation model designed for fully inductive KGC. We evaluate ULTRA under the 0-shot setting. See Appendix E for more details on the versions of ULTRA used. The results are in Figure 6(b) where for each dataset, we average the results across the different inference graphs (full results in Appendix A.2). On the (E) task, ULTRA is comparable to the supervised SOTA. However, on the (E, R) task, ULTRA significantly outperforms other methods. This suggests that ULTRA contains a much greater ability than other existing methods to generalize to unseen relations.
(a) PPR vs. SOTA Hits@10 on newer datasets
(b) Mean Hits@10 for ULTRA vs. SOTA
Figure 6: (a) Performance of PPR vs. SOTA on new inductive datasets. Note that we abbreviate CoDEx-M as CM, FB15k-237 as FB, HetioNet as HN and WN18RR as WN. Furthermore, the number represents the inference graph. (b) Mean performance of the foundation model ULTRA [34] vs. the supervised SOTA on the newer datasets.
6 Conclusion
@@
In this paper we study the problem of constructing datasets for inductive knowledge graph completion. Upon examination, we find that we can achieve competitive performance on most inductive datasets through the use of Personalized PageRank [12], which ignores the relational structure of the graph. Through our study, we uncover that this shortcut is due to how inductive datasets are created. To remedy this problem, we propose a new dataset construction process based on graph partitioning that empirically mitigates the impact of the studied shortcut. We then construct new benchmark datasets using this new procedure and benchmark various methods. Examining the results, we observe that the relative performance decreases on most datasets and methods, indicating that the newer benchmark datasets are harder than the previous ones. For future work, we plan to explore creating inductive KG datasets that aren’t sampled from existing transductive datasets.
@@ Appendix A Additional Results
A.1 Supervised SOTA vs. PPR Performance
In this section we show the performance in terms of Hits@10 for the SOTA supervised method vs. the Personalized PageRank (PPR) score. We further include the % difference in performance and which method is considered SOTA. These are shown in Tables 67, and 8 for the (E), (E, R), and transductive datasets, respectively. For an overview of SOTA performance on additional KG datasets, please see [34].
Table 6: Hits@10 for PPR vs. Supervised SOTA on the (E) Inductive Datasets
@@
WN v1 82.6 77.1 7% NBFNet [8]
WN v2 79.8 74.4 12% NBFNet [8]
WN v3 56.8 45.2 26% NBFNet [8]
WN v4 74.3 67.3 10% A*Net [16]
FB v1 60.7 41.2 47% NBFNet [8]
FB v2 70.4 47.6 48% NBFNet [18]
FB v3 66.7 43.5 53% NBFNet [8]
FB v4 66.8 38.4 74% NBFNet [8]
ILPC-S 25.1 19.8 27% NodePiece [19]
ILPC-L 14.6 22.5 -35% NodePiece [19]
Table 7: Hits@10 for PPR vs. Supervised SOTA on the (E, R) Inductive Datasets
@@
FB-100 37.1 22.2 67% InGram [9]
FB-75 32.5 21.9 48% InGram [9]
FB-50 21.8 20.5 6% InGram [9]
FB-25 27.1 20.9 30% InGram [9]
WK-100 16.9 15.8 7% InGram [9]
WK-75 36.2 29.5 23% InGram [9]
WK-50 13.5 10.6 27% InGram [9]
WK-25 30.9 23.2 33% InGram [9]
Table 8: Hits@10 for PPR vs. Supervised SOTA on the Transductive Datasets
@@
FB15k-237 66.6 2.7 2367% NBF+TAGNet [18]
WN18RR 59.9 46.2 30% NBF+TAGNet [18]
CoDEx-M 49.0 9.0 444% ComplEx RP [35]
CoDEx-S 66.3 8.6 671% ComplEx RP [35]
CoDEx-L 47.3 9.0 426% ComplEx RP [35]
Hetionet 40.3 2.4 1579% RotatE [36]
DBPedia100k 41.8 30.2 38% ComplEx-NNE+AER [37]
A.2 Performance of ULTRA
We include the full results of ULTRA [34] on each dataset and inference graph. The result on the (E) datasets are in Table 9 while those for the (E, R) datasets are in Table 10.
Table 9: (E) Inductive Results for ULTRA [34]
Metric CoDEx-M WN18RR HetioNet
Inference 1 Inference 1 Inference 2 Inference 1 Inference 2
MRR 30.2 64.7 21.4 57.9 72.7
Hits@10 46.6 72.7 38.5 69.1 86.3
Table 10: (E) Inductive Results for ULTRA [34]

### v3

@@ 3.3. Why is This So Common on Inductive Datasets?
Discussion. The studies in this section show that the current strategy for constructing inductive datasets from existing transductive datasets is liable to introduce a well-performing shortcut into the existing graph. Currently, it almost always leads to a sharp increase in the PPR performance. Furthermore, attempting to limit the severity of this issue is very difficult while also generating train and inference graphs of reasonable sizes. This suggests that a new strategy is needed for sampling graphs for inductive KGC from transductive datasets.
@@
Figure 4. Experiments when generating inductive datasets from FB15k-237. (a) The generated datasets always have a much higher $\Delta\text{SPD}$ than the transductive dataset. Furthermore, there is a trade-off between the size of the train (b) and inference graph (c), making it difficult to obtain a good split.
4. Constructing Inductive Datasets through Graph Partitioning
@@
4.1. Partition-Based Dataset Sampling
@@
We note that the task of constructing inductive datasets from an existing graph $\mathcal{G}$ can be framed as a graph partitioning problem. Formally, we want to sample two non-overlapping partitions from the graph such that $\mathcal{G}_{\text{train}},\mathcal{G}_{\text{inf}}\subseteq\mathcal{G}$ and $\mathcal{G}_{\text{train}}\cap\mathcal{G}_{\text{inf}}=\emptyset$. Given the analysis in Section 3.2, we hope to sample subgraphs such that $\Delta\text{SPD}(\mathcal{G}_{\text{train}})\approx\Delta\text{SPD}(\mathcal{G}_{\text{inf}})\approx\Delta\text{SPD}(\mathcal{G}$). But, how do we find subgraphs that satisfy this property? Intuitively, we want to sample each subgraph so that it’s removal has little effect on the initial graph’s structure. We give an example in Figure 5 where we sample two subgraphs from an existing graph. As we can see, even though the graph is partitioned, the relationship between entities in the same partition remain roughly the same before and after the partition. This is because there already exists little relationship between the two partitions in the original graph.
@@
Multiple popular approaches (Shi and Malik, 2000; Blondel et al., 2008) exist that attempt to divide the graph into optimal partitions. The guiding principle in these approaches is that the partitions should be internally dense but only sparsely connected to one another. Because of this, the entities in different partitions should only be weakly connected and have little impact on one another. Therefore, the relationship between entities in the same partition are minimally affected by outside entities or edges. As such, removing this partition from the graph should then have little effect on the entities in that partition. Also, since the partitions are created at the same time, we avoid sampling the inference graph after the training, which can be suboptimal.
Figure 5. Sampling Two Subgraphs
@@
In practice, we consider using Spectral Clustering (Shi and Malik, 2000) or the Louvain method (Blondel et al., 2008), as dependent on the dataset. Once the original graph is partitioned into $n$ partitions, we sample $k$ of those to be used. The partitions are chosen such that they display similar properties to the original graph (see Section 4.2 for more). Of the $k$ partitions, one is chosen as the training graph while the other $k-1$ are designated as separate inference graphs. This is an important advantage of using partitioning to sample the graphs, as it allows us to avoid having to sample multiple different train and inference pairs as in (Teru et al., 2020; Lee et al., 2023), thereby allowing for more efficient benchmarking. Lastly, we note that when sampling graphs for the (E) inductive task, we further remove new relations from the inference graph. In practice, we find that we can easily find partitions where this amounts to little or no change in the inference graph. See Section B.2 for more details on the dataset creation process.
@@
Table 2. Mean $\Delta\text{SPD}$ and PPR Hits@10 of new and old Inductive (“Ind.”) datasets vs. their Transductive (“Trans.”) parent. We highlight the value of the inductive dataset closer to the transductive as blue.
@@ 4.2. Analysis of New Datasets
In this section we analyze the new inductive datasets created following the partition-based procedure outlined in Section 4.1. We create (E) inductive datasets, from WN18RR (Dettmers et al., 2018), CoDEx-M (Safavi and Koutra, 2020), and HetioNet (Himmelstein et al., 2017). For the (E, R) datasets, we use CoDEx-M (Safavi and Koutra, 2020) and FB15k-237 (Toutanova and Chen, 2015). Note that some datasets are only suitable for one task or another. For example, WN18RR and HetioNet have very few relations, making it nearly impossible to sample two partitions with significantly different relations for the (E, R) task. On the other hand, FB15k-237 contains too many relations to sample multiple graphs for the (E) tasks without removing many edges from either graph.
@@
In Table 2 we show the PPR Hits@10 and $\Delta\text{SPD}$ for the inductive datasets and their original transductive dataset. When multiple inference graphs exist, we take the mean across each inference graph. When possible, we also include a comparison against those inductive datasets that already exist. For example, for WN18RR in the (E) task, 4 datasets exist from (Teru et al., 2020). Compared to the old inductive datasets, the PPR performance for the newer datasets is much lower. Specifically, the average PPR performance is 78% lower on the new inductive datasets as compared to the older datasets. Also, the PPR performance of the new inductive datasets is very similar to the performance on the original transductive dataset. A similar trend can be found when comparing the $\Delta\text{SPD}$. This analysis shows that newer sampling procedure can indeed sample inductive datasets that are much more similar to the original transductive graph, greatly mitigating the PPR shortcut.
5. Experiments
5.1. Experimental Settings
@@
Datasets: We use the new datasets created in Section 4.1. For the (E) setting, this includes WN18RR, HetioNet, and CoDEx-M. For the (E, R) setting, it is FB15k-237 and CoDEx-M. We sample two inference graphs for each dataset, except for CoDEx-M on the (E) setting where we could only find one suitable graph for inference. For each inference graph, 10% of edges are randomly removed for testing. For validation 10% of edges are removed from the training graph. Note that while previous studies (Teru et al., 2020; Galkin et al., 2021) validate against samples from the inference graph, it is necessary that the validation samples be extracted from the train graph. This is because using validation samples from the inference graph can be seen as a form of test leakage, as we are able to see the inference graph before testing. As such, because the inference graphs must remain unobserved during training, we can only sample validation from the training graph. The statistics for each dataset can be found in Appendix B.2.
Evaluation: For a given test triple $(s,r,o)$, we strive to predict both $s$ and $o$ individually. This is framed as a ranking problem where we want the probability of the correct entity (e.g., $o$) to rank higher than all possible negative entities. Given the rank of the true entity among the negatives, we calculate both the Hit@K and the mean reciprocal rank (MRR), using the filtered setting following (Bordes et al., 2013). Following previous work (Galkin et al., 2023; Zhang and Yao, 2022; Lee et al., 2023), we evaluate against all possible negative entities.
@@
Baseline Methods: We consider prominent KGC methods including NBFNet (Zhu et al., 2021), RED-GNN (Zhang and Yao, 2022), NodePiece (Galkin et al., 2021), InGram (Lee et al., 2023), DEq-InGram (Gao et al., 2023), and Neural LP (Yang et al., 2017). We also consider the recent foundation model ULTRA (Galkin et al., 2023). Note that we were unable to obtain results ISDEA+, as the current code only works properly when evaluating against 50 random negatives. We also omit methods that have been shown to either be prohibitively expensive (e.g.,  (Teru et al., 2020; Liu et al., 2021; Mai et al., 2021)) or require the use of textual information (Gesese et al., 2022; Daza et al., 2021).
Hyperparameters: Due to the different types and ranges of hyperparameters used for each method, the exact hyperparameters and their ranges differ by method. For each method, we tried to follow the recommended ranges used by the authors. For InGram (Lee et al., 2023) we tuned the learning rate in $\{1e^{-3},5e^{-4}\}$ and the number of entity layers in $\{2,4\}$. For NodePiece (Galkin et al., 2021) we tuned the learning rate from $\{1e^{-3},1e^{-4}\}$ and the loss margin from $\{15,25\}$. For RED-GNN (Zhang and Yao, 2022), we tuned the learning rate from $\{5e^{-3},5e^{-4}\}$ and the dropout from $\{0.1,0.3\}$. In their experiments, NBFNet (Zhu et al., 2021) uses the same hyperparameters across all datasets. We therefore do the same. For Neural LP (Yang et al., 2017) we also use the default set of hyperparameters that is shared across datasets.
Training Settings: Each model was trained on either a: Tesla V100 32Gb, NVIDIA RTX A6000 48Gb, NVIDIA RTX A5000 24Gb, or Quadro RTX 8000 48Gb. All models were implemented in PyTorch (Paszke et al., 2019) and Torch-Geometric (Fey and Lenssen, 2019).
5.2. Results
@@
Main Results: The main results for the supervised methods can be found in Tables 3 and 4. We observe that on nearly every dataset, NBFNet and RED-GNN are the two best models. This indicates that conditional MPNNs (Huang et al., 2024), the class of model in which both belong to, are necessary for strong performance in inductive KGC. Interestingly, we observe that InGram struggles in the (E, R) setting. This is even true when the % of new relations is high. This runs counter to the results on older inductive datasets (Lee et al., 2023) where InGram excelled over NBFNet and RED-GNN. However, this is not true for DEq-InGram, which performs consistently well under the (E, R) setting. We also observe that Neural LP sometimes fails to converge, resulting in a near zero performance and thus causing the model to have a high performance variance. Lastly, the results when evaluating via MRR can be found in Appendix A.3 and are consistent with those shown here.
Performance Comparison of PPR vs. SOTA: In our original analysis, we showed that for the older inductive datasets, the performance gap between the SOTA method of each dataset and PPR is quite small. Specifically, the mean % difference between the two was relatively small at 25-29% (see Figure 1). We now perform the same analysis on the new inductive datasets. The results are shown in Figure 6. We observe that PPR generally performs much worse than the SOTA method with a mean % difference of 101%. Furthermore, this is more in line with the transductive datasets which have a mean % difference of 121%. This suggests that in addition to the decrease in raw PPR performance, the relative performance of PPR also decreases on the newer datasets.
Table 3. (E) Inductive Results (Hits@10) for supervised methods.
@@
Table 4. (E, R) Inductive Results for supervised methods. The % of new relations are in parentheses.
@@
Figure 6. PPR vs. SOTA Hits@10 on new datasets.
Table 5. Mean performance on Old vs. New inductive datasets by Method. Red indicates a decrease in the mean performance on the newer splits.
@@
Performance Comparison on New and Old Inductive Datasets: In Table 2, we compared both the PPR performance and $\Delta\text{SPD}$ on the new and older inductive datasets. We found that both metrics tend to be much higher on the older inductive datasets, indicating that our newer splits are effective in mitigating the shortcut. Given those results, a natural question is whether we see a similar drop in performance for neural methods? We limit our analysis to WN18RR (E) and FB15k-237 (E, R). This is either due to a lack of older datasets (i.e., CoDEX-M (E, R) and HetioNet (E)) or minimal results on the older datasets, i.e., CoDEX-M (E). For each dataset, we compute the mean performance across inference graphs. The results are shown in Table 5. We find that performance of most methods drops significantly on both datasets, with an average drop of 40.6% and 9.5% on WN18RR (E) and FB15k-237 (E, R), respectively. This suggests that removing the shortcut has a large negative effect on the performance, suggesting that our new datasets are indeed harder.
@@
Performance of ULTRA (Galkin et al., 2023): We further compare against ULTRA, a recent foundation model designed for fully inductive KGC. We evaluate ULTRA under the 0-shot setting. Since the setting of ULTRA differs from that of the other methods (i.e., 0-shot), we display it separately from the other methods in Tables 3 and 4. See Appendix D for more details on the versions of ULTRA used. The results are in Figure 7 where for each dataset, we average the results across the different inference graphs (full results in Appendix A.2). On the (E) task, ULTRA is comparable to the supervised SOTA. However, on the (E, R) task, ULTRA significantly outperforms other methods. This suggests that ULTRA contains a greater generalization ability than other methods.
Figure 7. Mean Hits@10 for ULTRA vs. Supervised SOTA
6. Conclusion
@@
In this paper we study the problem of constructing datasets for inductive knowledge graph completion. Upon examination, we find that we can achieve competitive performance on most inductive datasets through the use of Personalized PageRank (Page et al., 1999), which ignores the relational structure of the graph. Through our study, we uncover that this shortcut is due to how inductive datasets are created. To remedy this problem, we propose a new dataset construction process based on graph partitioning that empirically mitigates the impact of the studied shortcut. We then construct new benchmark datasets using this new procedure and benchmark various methods. Examining the results, we observe that the relative performance decreases on most datasets and methods, indicating that the newer benchmark datasets are harder than the previous ones. For future work, we plan to explore creating inductive KG datasets that aren’t sampled from existing transductive datasets.
Acknowledgements.
This research is supported by the National Science Foundation (NSF) under grant numbers CNS2321416, IIS2212032, IIS2212144, IOS2107215, DUE2234015, CNS2246050, DRL2405483 and IOS2035472, US Department of Commerce, Gates Foundation, the Michigan Department of Agriculture and Rural Development, Amazon, Meta, and SNAP.
@@ Appendix A Additional Results
A.1. Supervised SOTA vs. PPR Performance
In this section we show the performance in terms of Hits@10 for the SOTA supervised method vs. the Personalized PageRank (PPR) score. We further include the % difference in performance and which method is considered SOTA. These are shown in Tables 67, and 8 for the (E), (E, R), and transductive datasets, respectively. For an overview of SOTA performance on additional KG datasets, please see (Galkin et al., 2023).
Table 6. Hits@10 for PPR vs. Supervised SOTA on the (E) Inductive Datasets
@@
WN v1 82.6 77.1 7% NBFNet (Zhu et al., 2021)
WN v2 79.8 74.4 12% NBFNet (Zhu et al., 2021)
WN v3 56.8 45.2 26% NBFNet (Zhu et al., 2021)
WN v4 74.3 67.3 10% A*Net (Zhu et al., 2023)
FB v1 60.7 41.2 47% NBFNet (Zhu et al., 2021)
FB v2 70.4 47.6 48% NBFNet (Shomer et al., 2023b)
FB v3 66.7 43.5 53% NBFNet (Zhu et al., 2021)
FB v4 66.8 38.4 74% NBFNet (Zhu et al., 2021)
ILPC-S 25.1 19.8 27% NodePiece (Galkin et al., 2021)
ILPC-L 14.6 22.5 -35% NodePiece (Galkin et al., 2021)
Table 7. Hits@10 for PPR vs. Supervised SOTA on the (E, R) Inductive Datasets
@@
FB-100 37.1 22.2 67% InGram (Lee et al., 2023)
FB-75 32.5 21.9 48% InGram (Lee et al., 2023)
FB-50 21.8 20.5 6% InGram (Lee et al., 2023)
FB-25 27.1 20.9 30% InGram (Lee et al., 2023)
WK-100 16.9 15.8 7% InGram (Lee et al., 2023)
WK-75 36.2 29.5 23% InGram (Lee et al., 2023)
WK-50 13.5 10.6 27% InGram (Lee et al., 2023)
WK-25 30.9 23.2 33% InGram (Lee et al., 2023)
Table 8. Hits@10 for PPR vs. Supervised SOTA on the Transductive Datasets
@@
FB15k-237 66.6 2.7 2367% NBF+TAGNet (Shomer et al., 2023b)
WN18RR 59.9 46.2 30% NBF+TAGNet (Shomer et al., 2023b)
CoDEx-M 49.0 9.0 444% ComplEx RP (Chen et al., 2021)
CoDEx-S 66.3 8.6 671% ComplEx RP (Chen et al., 2021)
CoDEx-L 47.3 9.0 426% ComplEx RP (Chen et al., 2021)
Hetionet 40.3 2.4 1579% RotatE (Sun et al., 2019)
DBPedia100k 41.8 30.2 38% ComplEx-NNE+AER (Ding et al., 2018)
A.2. Performance of ULTRA
We include the full results of ULTRA (Galkin et al., 2023) on each dataset and inference graph. The result on the (E) datasets are in Table 9 while those for the (E, R) datasets are in Table 10.
Table 9. (E) Inductive Results for ULTRA (Galkin et al., 2023)
Dataset Graph Hits@10 MRR
CoDEx-M Inference 1 30.2 46.6
WN18RR Inference 1 64.7 72.7
Inference 2 21.4 38.5
HetioNet Inference 1 57.9 69.1
Inference 2 72.7 86.3
Table 10. (E) Inductive Results for ULTRA (Galkin et al., 2023)